Find a quadratic function with the roots 1 and 3.
Don't have much experience with these. Any help is appreciated!

Found 2 solutions by josgarithmetic, addingup:
Answer by josgarithmetic(39613)
Answer by addingup(3677)

Let f(x) be the function where 1 and 3 are the roots. Then, the factors are (x-1) and (x-3)
f(x) = (x-1)(x-3)
Now FOIL to open the brackets (First Outer Inner Last) and you get:
f(x) = x^2-3x-1x+3 = x^2-4x+3
:
John